Tohban Report 2014-05-14
From RHESSI Wiki
|Start Date:||7 May 2014|
|End Date:||14 May 2014|
|Tohban:||Juan Carlos Martinez Oliveros (Vice Tohban: Albert Shih)|
|Next Tohban:||Lindsay Glesener (Vice Tohban: Sam Krucker)|
|List all reports|
Solar activity this week was low. There are currently 10 active regions on the disk. Farside monitoring shows an active region that may rotate into the disk in the following day.
How many GOES flares occurred?
Flares above B, C, M, X class were 1 36 2 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 1 26 1 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 1 13 0 0
There were RHESSI flares/GOES flares 205 / 39 over the time range 07-May-14 14-May-14
The SSR was not emptying by the end of the daily pass set, reaching a maximum of 41% and a minimum of 10%.
The decimation was changed on 2014 May 12 from normal/vigorous to active/vigorous.
The cold plate 1 and 2 are at 117.2 K and 115.8 K respectively.
No data gaps this week
Detector 8 rear lifetime sharply decreased on 2014 May 9-10. The detector HV was drop by 10 counts.
Detector 4 rear threshold was changed from 0x70 to 0x80
Detector 4 also presented a strange noise on 2014 May 10 at about 22:40.
Some anomalies were reported this week. First the state of health reports were saturated for 50 minutes due a ITOS system crash on Monday. The TLE system was stalled during the weekend, which caused dropouts during high elevations passes. All the data were replayed. The thin shutter was in between 2014 May 13 at 22:09 and 2014 May 14 at 01:11, this issue is being investigated.
No operations this week
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 minutes|
|Night time data (rears)||+/- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Detector problems?||4, 8|